OpenStudy (anonymous):

x + y = $39.00 x + 2y = $45.50

OpenStudy (anonymous):

elimination*

OpenStudy (rogue):

Subtract the first equation from the second to instantly get y.

OpenStudy (rogue):

Then plug in y into equation 1 to get x. Done. :)

OpenStudy (anonymous):

-1y = -6.5 ?

OpenStudy (rogue):

Yeah, so +y = +6.5. Then plug that back in to find x.

hero (hero):

x + y = $39.00 x + 2y = $45.50 x = 39 - y x = 45.5 - 2y x = x 39 - y = 45.5 - 2y 2y - y = 45.5 - 39 y = 6.5

OpenStudy (rogue):

Too long hero, its easier to subtract/add than substitute :P

hero (hero):

How come you didn't do it then? I could have skipped steps if I wanted, but for the sake of demonstration, included most steps.

OpenStudy (anonymous):

btw. for; 6x + 4y = 1128 would x = 108 and y = 120 ?

OpenStudy (rogue):

Didn't mean to offend you or anything, but I just said that doing the problem by subtracting the equations is a faster method. I didn't do it because its better that Justine gets practice doing it herself...

OpenStudy (rogue):

Are you given a second equation in addition to 6x + 4y = 1128.

OpenStudy (anonymous):

I made is from a word problem, but not sure. Hold on 1 min

OpenStudy (anonymous):

At all you can eat restaurant bbq fundraiser that you are sponsoring, adults pay $6 and kids pay $4 for a dinner. 212 people attend, and you raise $1128. How many adults attended? How many kids?

hero (hero):

Rogue, you can't offend me and I do things my way. No offense, but I don't need you to tell me what's easier. I prefer to use my own methods.

OpenStudy (rogue):

Alright, thats fine with me, Hero :) For the word problem, let x = # of parents, y = of kids. x + y = 212 6x + 4y =1128 Now just solve the system.